Sarasota offers an MP3 download (seven-single medley) as a premium for every 5 candy bar wrappers presented by customers together with $3.30. The candy bars are sold by the company to distributors for 30 cents each. The purchase price of each download code to the company is $3.05. In addition, it costs 50 cents to distribute each code. The results of the premium plan for the years 2020 and 2021 are as follows. (All purchases and sales are for cash.)

2020

2021

MP3 codes purchased 340,000 448,800
Candy bars sold 2,665,300 2,676,600
Wrappers redeemed 1,632,000 2,040,000
2020 wrappers expected to be redeemed in 2021 394,400
2021 wrappers expected to be redeemed in 2022 476,000

Part 1

Prepare the journal entries that should be made in 2020 and 2021 to record the transactions related to the premium plan of the Sarasota. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 0 decimal places, e.g. 1,525.)

Account Titles and Explanation

Debit

Credit

2020

(To record the premium inventory.)

(To record the sales.)

(To record the expense associated with the sale.)

(To record the premium liability.)

2021

(To record the premium inventory.)

(To record the sales.)

(To record the expense associated with the sale.)

(To record the premium liability.)

eTextbook and Media

List of Accounts

Save for Later

Attempts: 0 of 15 used

Submit Answer

Part 2

Indicate the amounts for each accounts, and classifications of the items related to the premium plan that would appear on the balance sheet and the income statement at the end of 2020 and 2021.

Amount

Account

2020

2021

Classification

Inventory of Premiums $ $                                                                       Long-term InvestmentsSelling ExpenseCurrent LiabilityCurrent AssetProperty, Plant and EquipmentStockholders' Equity
Premium Liability                                                                       Long-term InvestmentsCurrent AssetProperty, Plant and EquipmentSelling ExpenseCurrent LiabilityStockholders' Equity
Premium Expense
Answer #1

Part 1
2020

Account Titles Debit Credit
Premium Inventory $     1,037,000 =340000*3.05
       Accounts Payable / Cash $ 1,037,000
(To record the premium inventory)
Cash / Accounts Receivable $        799,590 =2665300*0.3
       Sales Revenue     $     799,590
(To record the sales)
Cash $        913,920 =1632000/5*(3.3-0.5)
Premium Expense $          81,600 =1632000/5*(3.05+0.5-3.3)
      Premium Inventory $     995,520 =1632000/5*3.05
(To record the expense associated with the sale)
Premium Expense $          19,720 =394400/5*(3.05+0.5-3.3)
      Premium Liability $       19,720
(To record premium liability)

2021

Account Titles Debit Credit
Premium Inventory $     1,368,840 =448800*3.05
       Accounts Payable / Cash $ 1,368,840
(To record the premium inventory)
Cash / Accounts Receivable $        802,980 =2676600*0.3
       Sales Revenue     $     802,980
(To record the sales)
Cash $     1,142,400 =2040000/5*(3.3-0.5)
Premium Expense $          82,280 =2040000/5*(3.05+0.5-3.3)-19720
Premium Liability $          19,720
      Premium Inventory $ 1,244,400 =2040000/5*3.05
(To record the expense associated with the sale)
Premium Expense $          23,800 =476000/5*(3.05+0.5-3.3)
      Premium Liability $       23,800
(To record premium liability)
Account 2020 2021
Inventory of Premiums $          41,480 $     165,920 Current Asset
Premium Liability $          19,720 $       23,800 Current Liability
Premium Expense $        101,320 $     106,080

If you have any query, kindly comment with your query and please mark thumbs up.